OpenStudy (anonymous):

Write the given expression in terms of x and y only. sin(sin^−1 (x) + cos^−1 (y))

OpenStudy (anonymous):

Using the addition formula I get: sin(sin^-1(x))cos(sin^-1(y)) + cos(sin^-1x))sin(^-1(y)) But I'm not really sure how to go from there.

OpenStudy (anonymous):

hmm, i think it might be a little mixed up. The sum formula is:\[\sin(a+b)=\sin a \cos b+ \sin b \cos a\]substituting:\[a=\sin ^{-1} x,b=\cos^{-1}y\]we get:\[\sin (\sin ^{-1}x)\cos(\cos^{-1}y)+\sin(\cos ^{-1}y)+\cos(\sin ^{-1}x)\]

OpenStudy (anonymous):

Right. I did switch it up, but how am I supposed to get things in terms of X and Y from there ?

OpenStudy (anonymous):

the two terms on the left simplify nicely:\[\sin (\sin ^{-1}x)=x, \cos( \cos^{-1}y)=y\]

OpenStudy (anonymous):

What about the second half of the problem.

OpenStudy (anonymous):

The ones on the right might require some right triangle drawing.

OpenStudy (anonymous):

Continuing after @joemath314159: \[ \sin (\sin ^{-1}x)\cos(\cos^{-1}y)+\sin(\cos ^{-1}y)+\cos(\sin ^{-1}x) \] \[\sin (\sin ^{-1}x) = x\] \[ \cos(\cos^{-1}y) = y\] \[ \sin(\cos ^{-1}y) = \sqrt{1-y^2} \] \[ \cos(\sin ^{-1}x = \sqrt{1-x^2} \] Now substitute.

OpenStudy (anonymous):

Ahhh, many thanks.